OpenStudy (anonymous):

sqrt3x - 5 = 4 A.7 B.3 C.1/3 D.49

OpenStudy (callisto):

\[\sqrt{3x-5}=4 ?\]

OpenStudy (campbell_st):

looks like A

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

3 x 7 - 5 = 16

OpenStudy (callisto):

square both sides and you'll get 3x - 5 =4^2 3x-5 = 16 3x =21 Can you get the answer from here?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Glad to hear :)